# Self-regularized entropy: How much do quantum black holes violate the no-hair theorem?

### v1

@@ IV Mode counting and canonical entropy
In this section we derive the entropy of a single massless scalar field within the $q$-metric, at Hawking temperature.
@@ IV.1 WKB radial spectrum
For each $(\ell,m)$ the radial wave‑number is $k_{r}(\rho)=\sqrt{\omega^{2}-V_{\ell}(\rho)}$. With $r_{*}^{\min}=0$ and $r_{*}^{\max}=L$ (Sec. III.2). In the WKB approximation, for fixed $(\ell,m)$ the counting function $N_{\ell}(\omega)$ is the radial mode number with frequency $\leq\omega:=\omega_{n\ell}$ is smooth and given by
$n:=N_{\ell}(\omega)=\frac{1}{\pi}\int_{0}^{L}\,k_{r}\,dr_{*},$ (29)
inside the cavity we have $k_{r}\simeq\omega$, therefore $n\simeq\frac{\omega L(\omega,\ell)}{\pi}$ and from equations (25) and (26) we have
$L(\omega,\ell)=\frac{4m}{\beta}\,\left(\frac{4m^{2}\omega^{2}}{\ell^{2}}\right)^{\!\beta/(\alpha-2q)}.$ (30)
Writing $L(\omega,\ell)=C(q,m)\,\ell^{-s}\omega^{s}$ gives the level law $\omega:=\omega_{n\ell}\ \propto\ n^{\delta}\,\ell^{\sigma}$ with $\delta:=\frac{1}{1+s},\sigma:=\frac{s}{1+s}$, with
$\quad s:=\frac{2}{p}=\frac{2\beta}{\alpha-2q}=\frac{q^{2}}{1-q^{2}-q^{3}}>0,$ (31)
and
$C(q,m)=\frac{4m}{\beta}\,(4m^{2})^{\,\beta/(\alpha-2q)}=\frac{4m}{\beta}\,(4m^{2})^{\,s/2}.$ (32)
@@ IV.2 Radial density of states
By defining the dimensionless frequency $w=\omega/T_{H}$, the canonical entropy of one bosonic oscillator is $\mathcal{S}(w)=\dfrac{w}{e^{w}-1}-\ln\,\bigl{(}1-e^{-w}\bigr{)}$. The total entropy is then given by a discrete sum over all the modes:
$S(q)$ $\equiv\sum^{\infty}_{\ell=0}(2\ell+1)\sum_{n=1}^{\infty}\mathcal{S}\left(\frac{\omega_{n\ell}}{T_{H}}\right)\simeq\int_{0}^{\infty}2\ell d\ell\sum_{n=1}^{\infty}\mathcal{S}(w)$
@@
with $n=\left(\frac{C}{\pi}\right)\ell^{-s}\omega^{s+1}$, and differentiate at fixed $n$ finds
$\sum_{n=1}^{\infty}2\ell\,\frac{\partial\ell}{\partial w}=\frac{2(1+s)}{s}\,\left(\frac{C}{\pi}\right)^{\frac{2}{s}}\,\zeta\,\left(\frac{2}{s}\right)\,T_{H}^{\frac{2(1+s)}{s}}w^{\frac{2(1+s)}{s}-1}.$ (34)
@@
$S(q)=\frac{2(1+s)}{s}\,\left(\frac{C}{\pi}\right)^{\frac{2}{s}}\,\zeta\,\left(\frac{2}{s}\right)\,T_{H}^{\frac{2(1+s)}{s}}\int_{0}^{\infty}w^{\frac{2}{s}+1}\,\mathcal{S}(w)\,dw,$
@@
$S(q)=(p+2)\,\left(\frac{C}{\pi}\right)^{p}\,T_{H}^{p+2}\,\zeta(p)(p+3)\,\Gamma(p+2)\,\zeta(p+3).$
From the turning point analysis equation (26) and equation (30), and recall that $L(\omega,\ell)=C(q,m)\,\ell^{-s}\omega^{s}$ we can write
$\frac{C}{\pi}\,=\frac{4m}{\beta\pi}(4m^{2})^{s/2},\,$ (35)
using $T_{H}=(8\pi m)^{-1}$ and $p=\frac{2}{s}$ we find
$\left(\frac{C}{\pi}\right)^{p}\,T_{H}^{p+2}=(4m^{2})\left(\frac{4m}{\beta\pi}\right)^{p}(8\pi m)^{-(p+2)}$ $=2^{-(p+4)}\,\beta^{-p}\,\pi^{-(2p+2)},\,$
the factors of $m$ canceled out. Thus, the final result is
$S(q)=\frac{(p+2)(p+3)}{2^{\,p+4}\,\pi^{\,2p+2}}\;\beta^{-p}\;\Gamma(p+2)\,\zeta(p)\,\zeta(p+3),$ (36)
where $p\approx\frac{2}{q^{2}}$ and $\beta\approx q^{2}$ from equations (20) and (4). This entropy is finite for all sufficiently small $|q|\neq 0$. In contrast, as $q\to 0$ one has $p\to\infty$ (hence $s(q)\to 0^{+}$), the cavity length reverts to the Schwarzschild logarithm, and the entropy reproduces the usual brick wall divergence. In the formal zeta-regularized notation used earlier, this appears as the symbolic limit $\bigl{[}2^{-3}/q^{4}\bigr{]}\Gamma(0)\,\zeta(1)$ when $q\to 0^{+}$.
@@ V Observational prospects
For any four‑dimensional, non‑extremal black hole
@@
where $m_{\mathrm{P}}=2.18\times 10^{-5}\,\mathrm{g}$, is the Planck mass  [^2] For nonextremal Kerr, $8\pi M^{2}\leq A_{H}\leq 16\pi M^{2}$ in $G=c=\hbar=1$, so $\ln S_{\rm BH}$ changes by at most $\ln 2$, negligible for our purposes.. Therefore, $\ln S_{\rm BH}$ grows only logarithmically with the mass. Using the bound (3) we set earlier in Section V.1,
@@
we find that across the entire stellar to supermassive interval $5\,M_{\odot}\leq M\leq 10^{10}\,M_{\odot}$, leading to $S_{\rm BH}\,\sim\,10^{78}\text{--}10^{97}$, the bound varies only from $0.095$ to $0.11$ (light stellar) to $0.095$ (most massive).
However, from the entropy that we derived in the previous section, equation (36), we can put a stronger bound on $q$. For the astrophysically relevant small $|q|$ (and hence large $p$), by applying $\zeta(p+k)\to 1$, and Stirling’s formula for $\Gamma(p+2)$, thus $S(q)$ reduces to $p\log p$ relation which we invert with the product-log (Lambert) function $W$, defined by $W(z)\,e^{W(z)}=z$ for $z>0$, and we obtain
$q(S)\approx\sqrt{\frac{4W\left(\frac{\ln S}{4\pi\sqrt{e}}\right)}{\ln S}}.$ (39)
Plugging $S_{BH}\simeq 10^{78}-10^{97}$, i.e., $\ln S\simeq 180-223$ into the estimated formula above we obtain the more accurate bound on $q$ as
@@
which varies by less than 5% across 10 orders of magnitude in mass, for the black hole entropies across the stellar to supermassive mass range. This weak dependence on entropy is a strength: it means that $10-20\%$ quantum corrections to Kerr geometry are essentially mass independent and set by the absolute value of $|q|\sim 0.2$, while the spacetime remains close to Schwarzschild beyond a few gravitational radii.
@@ V.1 From static to stationary: Multipoles
The analysis in this paper has deliberately focused on the static $q$-metric, which is the only input needed to establish the near horizon self-regularisation of the canonical entropy. However, in this section we briefly connect that static description to stationary (spinning) black hole observables, solely to indicate how a nonzero $q$ would manifest as controlled deviations from the Kerr multipole hierarchy; here we use it only as a phenomenological map from $q$ to measurable multipoles. The $q$-metric admits a stationary (spinning) extension in which the mass and current multipole moments $(M_{\ell},J_{\ell})$ depart perturbatively from the Kerr values $M_{\ell}+iJ_{\ell}=M(ia)^{\ell}$. To lowest orders one finds (e.g., Toktarbay_2014 ; Faraji_2022 ), namely we have
$M_{0}=m+q\sigma,$
@@
for the lowest mass and angular momentum multipoles, where $\sigma=\sqrt{m^{2}-a^{2}}$ (e.g., Toktarbay_2014 ; Faraji_2022 ).
@@
where $a_{*}\equiv a/m$ is the dimensionless spin parameter, which ranges between 1-1 and $1$ for Kerr geometry (with $a_{*}=0$ for Schwarzschild). We see that violations of no-hair theorem are encoded in $(Q_{1},Q_{2})$, thus probing the potential quantum corrections induced by a finite value $q$.
But what value of $q$ should we use for a spinning quantum black hole? While our derivation of equation (39), based on the entropy of Hawking radiation was for a static q-metric, the fact that the dimensionless $q$-parameter changes by $5\%$ across 10 orders of magnitude in mass, suggests that it may also be insensitive to BH spin. More conservatively, given that $a_{*}\to 0$ is a continuous limit of Kerr spacetime, we may speculate that $a_{*}\sim{\cal O}(1)$ would only yield ${\cal O}(1)$ changes in $q$, thus keeping it within 102010-20% range (at least away from the extremal limit).
Figure 1: (Top, a) Predicted dependence of minimum $q$ parameter on the mass and spin of astrophysical black holes (equation 39). (Bottom, b) The range of quantum corrections induced in the multipolar structure of the Kerr metric, for the same span of black hole masses/spins.
With this preamble, equation (39) will be used to forecast anticipated departures from Kerr geometry, due to quantum violations of the no-hair theorem, within the context of $(Q_{1},Q_{2})$ parameters. However, it is important to note that the precise dependence on spin should not be taken too seriously. Fig. (1a) shows the values of $|q|$ predicted from equation (39) for astrophysical black hole masses, while Fig. (1b) plots the resulting $Q_{1}$ and $Q_{2}$ parameters. For astrophysical entropies $S\,\sim\,10^{78}\,-\,10^{97}$ the microstate bound implies $q\simeq 0.2$, well within the regime where the series converges and small enough that the metric still approximates Schwarzschild outside a few gravitational radii; we shall come back to this point below.
@@ V.2 Observational signatures
With these predictions at hand, we shall next explore how two forthcoming astrophysical experiments can probe these quantum violations of the no-hair theorem.
@@ V.2.1 Next Generation Event Horizon Telescope (ngEHT)
At (sub)millimetre wavelengths, the brightness distribution around a supermassive black hole encodes strong lensing geometry at radii $r\sim\mathcal{O}(m)$. The most robust geometric observables are: (i) the photon ring diameter and its departures from circularity (e.g., axial ratio, quadrupolar distortions) and (ii) the image centroid (or ring centre) relative to the dynamical centre. These quantities depend primarily on the spacetime lensing map and are comparatively insensitive to the detailed emissivity once multifrequency coverage is used.
@@ V.2.2 Laser Interferometer Space Antenna (LISA)
Let us now turn our focus to potential observations of extreme-mass-ratio inspirals (EMRIs) by the space-based Laser Interferometer Space Antenna (LISA). In an EMRI, a compact object of mass $\mu\,\ll\,m$ orbits a massive black hole and slowly inspirals through $\sim 10^{4}-10^{5}$ cycles in band e.g., PhysRevD.95.103012 . The phasing and precession structure of the gravitational waveform encode the multipole structure of the central object. In Kerr geometry, one has $Q_{2}=-Ma^{2}$; any deviation of the mass quadrupole $Q_{2}$ from this relation accumulates a measurable phase shift over the radiation-reaction time. Of course, spin and quadrupole both affect precession, so one needs at least one additional independent multipolar handle to break the quadrupole-spin degeneracy, practically, this comes from combining multiple observables: the two precessions and the frequency‑dependent sideband structure (inclination/polar motion harmonics) and late inspiral behaviour near the last stable orbit. In our language, this is equivalent to measuring one extra Kerr-null combination beyond the quadrupole, so that the fit cannot trade spin against $q$. With LISA‑class signal‑to‑noise, long coherent baselines, and waveform models that include precession and higher harmonics, a distortion as large as $q\sim 0.2$ would be potentially identifiable PhysRevD.52.5707 ; PhysRevD.75.042003 ; PhysRevD.95.103012 ; 2023LRR….26….2A .

### v2

@@ IV Mode counting and canonical entropy
In this section we derive the canonical entropy of a single massless scalar field from WKB counting of the near-horizon cavity modes, evaluated at the asymptotic temperature $T_{\rm H}$ adopted in Sec. II.1.
@@ IV.1 WKB radial spectrum
For each $(\ell,m)$ the radial wave-number is $k_{r}(\rho)=\sqrt{\omega^{2}-V_{\ell}(\rho)}$. Let $r_{*}^{\min}(\omega)$ and $r_{*}^{\max}(\omega,\ell)$ denote the classical turning points (Sec. III.2). In the WKB approximation, for fixed $(\ell,m)$ the radial overtone number $n:=N_{\ell}(\omega)$ is given by
$n=\frac{1}{\pi}\int_{r_{*}^{\min}}^{r_{*}^{\max}}k_{r}\,dr_{*}\,.$ (29)
In the high-frequency regime relevant for the ultraviolet mode count, $k_{r}\simeq\omega$ throughout most of the cavity and the turning-point regions contribute only subleading $O(1)$ phase shifts. Moreover, $r_{*}^{\min}$ produces an $\omega$-independent shift in $n$ (Sec. III.2), so for the purpose of the mode density we may approximate $n\simeq\omega L(\omega,\ell)/\pi$ with $L(\omega,\ell)\simeq r_{*}^{\max}$. Using equation (25) we obtain
$L(\omega,\ell)\simeq\frac{4m}{\beta}\,\left(\frac{4m^{2}\omega^{2}}{\ell^{2}}\right)^{\,\beta/(\alpha-2q)}.$ (30)
Finally, writing $L(\omega,\ell)=C(q,m)\,\ell^{-s}\omega^{s}$ gives an expression for the mode frequency $\omega_{n\ell}$ in terms of the angular harmonic and overtone numbers, $\ell$ and $n$:
$\omega_{n\ell}=\left(\pi\over C\right)^{\frac{1}{1+s}}n^{\frac{1}{1+s}}\ell^{\frac{s}{1+s}},$ (31)
$s:=\frac{2}{p}=\frac{2\beta}{\alpha-2q}=\frac{q^{2}}{1-q^{2}-q^{3}},$ (32)
$C(q,m):=\frac{4m}{\beta}\,(4m^{2})^{\,\beta/(\alpha-2q)}=\frac{2(1+q^{2}+q^{3})}{q^{2}}(4m^{2})^{\frac{1+s}{2}}.$
Note that $s>0$ as long as $q<0.75$.
@@ IV.2 Radial density of states
Equipped with the spectrum of the near-horizon trapped modes, we can compute their thermal entropy.
The canonical entropy of a bosonic harmonic oscillator with frequency $\omega$ at temperature $T$ is given by
$\mathcal{S}(w)=\dfrac{w}{e^{w}-1}-\ln\,\bigl(1-e^{-w}\bigr),$ (34)
in terms of its dimensionless frequency $w:=\omega/T$.
Using the large-$\ell$ angular-density approximation discussed in Sec. III.1, we treat the sum over angular modes by the standard replacement $\sum_{m}\to(2\ell+1)$ and convert the $\ell$ sum to an integral, which is sufficient for the ultraviolet scaling. The total thermal entropy is then given by a discrete sum over all the modes:
$S(q,T)$ $\equiv\sum^{\infty}_{\ell=0}(2\ell+1)\sum_{n=1}^{\infty}\mathcal{S}\left(\frac{\omega_{n\ell}}{T}\right)\simeq\int_{0}^{\infty}2\ell d\ell\sum_{n=1}^{\infty}\mathcal{S}(w)$
@@
with $n=\left(\frac{C}{\pi}\right)\ell^{-s}\omega^{s+1}$, and differentiating at fixed $n$ gives
$\sum_{n=1}^{\infty}2\ell\,\frac{\partial\ell}{\partial w}=\frac{2(1+s)}{s}\,\left(\frac{C}{\pi}\right)^{\frac{2}{s}}\,\zeta\,\left(\frac{2}{s}\right)\,T^{\frac{2(1+s)}{s}}w^{\frac{2(1+s)}{s}-1}.$ (36)
@@
$S(q,T)=\frac{2(1+s)}{s}\,\left(\frac{C}{\pi}\right)^{\frac{2}{s}}\,\zeta\,\left(\frac{2}{s}\right)\,T^{\frac{2(1+s)}{s}}\int_{0}^{\infty}w^{\frac{2}{s}+1}\,\mathcal{S}(w)\,dw,$
@@
$S(q,T)=(p+2)\,\left(\frac{C}{\pi}\right)^{p}\,T^{p+2}\,\zeta(p)(p+3)\,\Gamma(p+2)\,\zeta(p+3).$
If we adopt the asymptotic Hawking temperature $T_{\rm H}=(8\pi m)^{-1}$ (Sec. II.1), and $p=\frac{2}{s}$ we find
$\left(\frac{C}{\pi}\right)^{p}\,T_{\rm H}^{p+2}=(4m^{2})\left(\frac{4m}{\beta\pi}\right)^{p}(8\pi m)^{-(p+2)}$ $=2^{-(p+4)}\,\beta^{-p}\,\pi^{-(2p+2)},\,$
the factors of $m$ cancel out. Thus, the final result is
$S(q,T)=\frac{(p+2)(p+3)}{2^{\,p+4}\,\pi^{\,2p+2}}\;\beta^{-p}\;\Gamma(p+2)\,\zeta(p)\,\zeta(p+3)\left(T\over T_{\rm H}\right)^{p+2},$ (37)
where $p\approx\frac{2}{q^{2}}$ and $\beta\approx q^{2}$ from equations (20) and (4). This entropy is finite for all sufficiently small $|q|\neq 0$. In contrast, as $q\to 0$ one has $p\to\infty$ (hence $s(q)\to 0^{+}$), the cavity length reverts to the Schwarzschild logarithm, and the entropy reproduces the usual brick wall divergence. In the formal zeta-regularized notation used earlier, this appears as the symbolic limit $\bigl[2^{-3}/q^{4}\bigr]\Gamma(0)\,\zeta(1)$ when $q\to 0$. In this limit, by using $\zeta(p+k)\to 1$, and Stirling’s formula for $\Gamma(p+2)$ we find:
$\ln S(q,T)=-\frac{2}{q^{2}}\ln\left(e\pi^{2}q^{4}\times\frac{T}{T_{\rm H}}\right)+{\cal O}\left(1\over q\right)$ (38)
@@ V Observational prospects
In what follows, we use $q$ purely as a phenomenological deformation parameter in the external multipole structure and lensing geometry, in the same spirit as other bumpy-metric parametrisations used in tests of the Kerr hypothesis; the forecasts are sensitivity estimates, not evidence that the exact static q-metric is realized in nature. For any four-dimensional, non-extremal black hole
@@
where $m_{\mathrm{P}}=2.18\times 10^{-5}\,\mathrm{g}$, is the Planck mass  [^3] For nonextremal Kerr, $8\pi M^{2}\leq A_{H}\leq 16\pi M^{2}$ in $G=c=\hbar=1$, so $\ln S_{\rm BH}$ changes by at most $\ln 2$, negligible for our purposes.. Therefore, $\ln S_{\rm BH}$ grows only logarithmically with the mass. Using the echo-time bound (3) obtained in Sec. 2,
@@
we find that across the stellar-to-supermassive interval $5\,M_{\odot}\leq m\leq 10^{10}\,M_{\odot}$, corresponding to $S_{\rm BH}\sim 10^{78}\text{--}10^{97}$, the bound varies only weakly, from $\sim 0.11$ at the stellar-mass end down to $\sim 0.095$ at the most massive.
However, using the closed-form canonical entropy equation (38), we can sharpen this estimate by matching $S(q,T)$ to the microscopic entropy scale $S_{\rm BH}$ and solving for the corresponding $q(S_{\rm BH})$ in the small-$|q|$ regime. For the astrophysically relevant small $|q|$ we can invert (38) with the product-log (Lambert) function $W$, defined by $W(z)\,e^{W(z)}=z$ for $z>0$ to obtain
$q(S)\approx\sqrt{\frac{4W\left(\frac{\ln S_{\rm BH}\times\sqrt{T/T_{\rm H}}}{4\pi\sqrt{e}}\right)}{\ln S_{\rm BH}}}.$ (41)
Plugging $T\approx T_{\rm H}$ and $S_{\rm BH}\simeq 10^{78}-10^{97}$ (i.e., $\ln S\simeq 180-223$) into the estimated formula above, we obtain the corresponding estimate
@@
which varies by less than 5% across 10 orders of magnitude in mass, for the black hole entropies across the stellar to supermassive mass range. Similarly, changing the ratio of black hole temperature, $T$ to the standard Hawking temeparture $T_{\rm H}$ by up to a factor of 2, which may be expected due to quantum effects near the horizon (e.g., Oshita et al. (2020a)), only changes $q$ by less than 6%. This weak dependence on entropy is useful: within this framework, the entropy-motivated deformation scale is essentially mass and temperature independent across the stellar-to-supermassive range, with $|q|\sim 0.2$. When translated (Sec. V.1) into stationary multipole diagnostics, this corresponds to percent-to-tens-of-percent departures from the Kerr multipole relations, while the exterior remains close to Schwarzschild beyond a few gravitational radii in the sense of a small quadrupolar deformation.
@@ V.1 From static to stationary: Multipoles
The analysis in this paper has deliberately focused on the static $q$-metric, which is the only input needed to establish the near horizon self-regularisation of the canonical entropy. However, in this section we briefly connect that static description to stationary (spinning) black hole observables, solely to indicate how a nonzero $q$ would manifest as controlled deviations from the Kerr multipole hierarchy; here we use it only as a phenomenological map from $q$ to measurable multipoles. The $q$-metric admits a stationary (spinning) extension in which the mass and current multipole moments $(M_{\ell},J_{\ell})$ depart perturbatively from the Kerr values $M_{\ell}+iJ_{\ell}=M(ia)^{\ell}$. To lowest orders one finds (e.g., Toktarbay and Quevedo (2014); Faraji et al. (2022)), namely we have
$M_{0}=m+q\sigma,$ (43)
@@
for the lowest mass and angular momentum multipoles, where $\sigma=\sqrt{m^{2}-a^{2}}$ (e.g., Toktarbay and Quevedo (2014); Faraji et al. (2022)).
@@
where $a_{*}\equiv a/m$ is the dimensionless spin parameter, which ranges between 1-1 and $1$ for Kerr geometry (with $a_{*}=0$ for Schwarzschild). We see that violations of the Kerr multipole relations are encoded in $(Q_{1},Q_{2})$: these Kerr-null combinations vanish for Kerr and acquire a controlled $O(|q|)$ signal in the stationary extension.
What value of $q$ should one use at nonzero spin? Our entropy calculation leading to equation (41) is based on the static near-horizon geometry and does not determine a first-principles $q(a_{*})$ relation. For the purpose of forecasting the size of Kerr-null multipole diagnostics, we therefore adopt (41) as an entropy-motivated fiducial scale and evaluate it using the macroscopic Bekenstein-Hawking entropy of a Kerr black hole with the same mass and spin, $S_{\rm BH}(m,a_{*})$. Since $\ln S_{\rm BH}$ varies only by $O(1)$ (at most $\ln 2$) between Schwarzschild and extremal Kerr at fixed mass, the inferred $|q|$ changes only weakly with spin in this benchmark procedure (Fig. 1a).
Figure 1: (Top, a) Entropy-motivated estimate of $|q|$ across astrophysical masses obtained from (41), shown for two macroscopic entropy benchmarks: Schwarzschild ($a_{*}=0$) and extremal Kerr ($|a_{*}|=1$), which differ only by an $O(1)$ factor in $S_{\rm BH}$ at fixed mass. (Bottom, b) The corresponding magnitudes of the Kerr-null multipole diagnostics $(Q_{1},Q_{2})$ in (44) as functions of spin, for the same $|q|$ scale.
With this preamble, we use equation (41) only to set an entropy-motivated scale for $|q|$ and translate it into the Kerr-null multipole diagnostics $(Q_{1},Q_{2})$ through (44). The spin dependence shown in Fig. 1a arises solely from the choice of macroscopic entropy benchmark $S_{\rm BH}(m,a_{*})$ and should not be interpreted as a derived prediction for $q(a_{*})$ in a spinning near-horizon geometry. Fig. 1b then shows the corresponding $Q_{1}$ and $Q_{2}$ signals. For astrophysical entropies $S_{\rm BH}\sim 10^{78}\text{--}10^{97}$ this procedure yields $|q|\sim 0.2$, within the small-$|q|$ regime used throughout and small enough that the exterior remains close to Schwarzschild outside a few gravitational radii; we return to this below.
@@ V.2 Observational signatures
With these predictions in hand, we now outline how two forthcoming experiments could constrain such departures from the Kerr multipole relations.
@@ V.2.1 Next Generation Event Horizon Telescope (ngEHT)
At (sub)millimetre wavelengths, the brightness distribution around a supermassive black hole encodes strong-lensing geometry at radii $r\sim\mathcal{O}(m)$. The most robust geometric observables are: (i) the photon-ring diameter and its departures from circularity (e.g., axial ratio and quadrupolar distortions) and (ii) the image centroid (ring centre) relative to the dynamical centre. These quantities depend primarily on the spacetime lensing map and are comparatively less sensitive to emissivity once multifrequency coverage and multi-epoch averaging are used.
Quantifying the resulting photon-ring distortions requires ray tracing in the full metric (or its stationary extension); here we restrict to order-of-magnitude sensitivity targets.
[…]
@@ V.2.2 Laser Interferometer Space Antenna (LISA)
Let us now turn our focus to potential observations of extreme-mass-ratio inspirals (EMRIs) by the space-based Laser Interferometer Space Antenna (LISA). In an EMRI, a compact object of mass $\mu\,\ll\,m$ orbits a massive black hole and slowly inspirals through $\sim 10^{4}-10^{5}$ cycles in band e.g., Babak et al. (2017). The phasing and precession structure of the gravitational waveform encode the multipole structure of the central object. In Kerr geometry, one has $M_{2}=-Ma^{2}$; any deviation of the mass quadrupole moment $M_{2}$ from this relation accumulates a measurable phase shift over the radiation-reaction time. Of course, spin and quadrupole both affect precession, so one needs at least one additional independent multipolar handle to break the quadrupole-spin degeneracy, practically, this comes from combining multiple observables: the two precessions and the frequency‑dependent sideband structure (inclination/polar motion harmonics) and late inspiral behaviour near the last stable orbit. In our language, this is equivalent to measuring one extra Kerr-null combination beyond the quadrupole, so that the fit cannot trade spin against $q$. With LISA-class signal-to-noise, multi-year coherent baselines, and waveform models incorporating precession and higher harmonics, deviations from the Kerr quadrupole can be constrained at the $10^{-4}$ level Ryan (1995); Barack and Cutler (2007); Babak et al. (2017); Amaro-Seoane et al. (2023). Consequently, an $\mathcal{O}(10\%)$ deviation in the relevant multipolar structure would be large compared to the projected statistical precision in idealized EMRI analyses, and could be detectable with LISA provided waveform systematics and parameter degeneracies are under control.
V.2.3 Ground-Based Gravitational Wave Observatories
We can also compare our predictions with current bounds from LIGO-Virgo-KAGRA (LVK) ground-based gravitational wave observatories. To connect with LVK-style parametrisations, one often introduces a spin-induced quadrupole coefficient, $\kappa$, via $M_{2}=-\kappa\,\chi^{2}M_{0}^{3}$, where $\chi:=J_{1}/M_{0}^{2}$. From our definition of $Q_{1}$ in (44), one has the exact identity
$Q_{1}=\chi^{2}\,|\kappa-1|\qquad\Rightarrow\qquad|\kappa-1|=\frac{Q_{1}}{\chi^{2}},$ (46)
Combining (44) with $Q_{1}$ in equation (46) and using $\chi\simeq|a_{*}|$ at leading order in small $|q|$ yields the approximate translation
$|q|\simeq\frac{3}{2}\,\frac{a_{*}^{2}}{(1-a_{*}^{2})^{3/2}}\,|\kappa-1|\,$, (47)
[…]